x^-2-64y^-2
___________
9y-72x
i'm having trouble with this one.

Answer by sdmmadam@yahoo.com(530)

x^-2-64y^-2
___________
9y-72x
=[(1/x^2)-(64/y^2)]/(9y-72x) [using a^-m = (1/a^m)]
=(y^2-64x^2)/[9x^2y^2(y-8x)]
[finding the lcm in the nr and taking out 9 in the dr]
=[(y+8x)(y-8x)]/[9x^2y^2(y-8x)]
[using formula p^2-q^2 = (p+q)(p-q) and here p =y and q = 8x]
=(y+8x)/9x^2y^2 (cancelling (y-8x)in the nr and in the dr)